3x + 2y = 16
2x + y = 9

Answer by fractalier(6550)

From
3x + 2y = 16
2x + y = 9
solve the second equation for y and substitute it into the first equation...
y = -2x + 9
then
3x + 2(-2x + 9) = 16
3x - 4x + 18 = 16
-x = -2
x = 2
then
y = -2(2) + 9
y = 5